OpenStudy (anonymous):

Can someone help me work out this word problem? Essay: Show all work. A school designer wants to create a whiteboard with the optimal dimensions to enhance learning. It is determined that if one side of the whiteboard is x + 2y inches, the other side should be x^2+3xy-4y^2 inches. Write an algebraic expression for the area of such a whiteboard, simplify it, and include correct units with your solution.

OpenStudy (anonymous):

OK, area would be side times side, so you multiply the two expressions: (x + 2y in)(x^2 + 3xy - 4y^2 in) = x^3 + 3x^2y - 4xy^2 + 2x^2y + 6xy^2 - 8y^3 in^2 = x^3 + 5x^2y + 2xy^2 - 8y^3 in^2